How many liters of a 70% methyl solution should be added to 200 liters of a 20% methyl solution to obtain a 35% methyl solution?

Answer by ptaylor(2198)

Let x=amount of 70% methyl solution needed
Now we know that the amount of pure methyl that exists before the mixture takes place(0.70x+0.20*200) has to equal the amount of pure methyl that exists after the mixture takes place (0.35(200+x)). Sooooo:
0.70x+0.20*200=0.35(200+x) simplify
0.70x+40=70+0.35x subtract 40 and also 0.35x from each side
0.70x-0.35x=70-40 collect like terms
0.35x=30
x=85.7 liters amount of 70% methyl needed
CK
0.70*85.7+0.20*200=0.35*285.7
59.99+40=99.995
99.99~~~~99.995
Hope this helps---ptaylor